How do you solve
17- c = 19

Found 3 solutions by MathLover1, Alan3354, ikleyn:
Answer by MathLover1(20849)
Answer by Alan3354(69443)

How do you solve
17- c = 19
---
Subtract 17
-c = 2
Mutliply by -1
c = -2

Answer by ikleyn(52781)

.
Step 1.  Add " c " to both sides. You will get


         17 - c + c = 19 + c,  or   17 = 19 + c.



Step 2.  Subtract 19 from both sides.  You will get


         17 - 19 = 19 + c - 19,   or  -2 = c.


It is your answer:  c = -2.

Solved, explained and completed.

------------------

To learn this subject, read my lesson
    - HOW TO solve a linear equation
in this site.


Happy learning !